OpenStudy (anonymous):

Shawn and Ellie are construction workers. Shawn can complete a concrete job in 4 hours, while Ellie can complete it in 6 hours. The foreman says that it will take them 5 hours to complete it if they work together. Explain each step in solving this equation, and determine if the foreman is correct or not.

OpenStudy (kropot72):

If Shawn works for one hour he will complete 1/4 of the job. If Ellie works for one hour she will complete 1/6 of the job. If they work together for an hour they will complete 1/4 + 1/6 of the job. \[\large \frac{1}{4}+\frac{1}{6}=\frac{5}{12}\ ..........(1)\] So, working together for one hour, they will complete 5/12 of the job. Let x be the number of hours working together need to complete the job. Then the following equation can be written: \[\large \frac{5x}{12}=1\ ...............(2)\] Now you just need to solve equation (2) to find the time needed to complete the job when the two people work together.

OpenStudy (anonymous):

would it be x=12/5

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Good work! You are correct. \[\large \frac{12}{5}=2.4\ hours\ or\ 2\ hours\ 24\ minutes\]

OpenStudy (anonymous):

cool thanks @kropot72

OpenStudy (kropot72):

You're welcome :)